OpenStudy (anonymous):

It costs $350 to carpet an 8-ft-by-10-ft room. At this rate, how many dollars would it cost to put the same carpeting in a 24-ft-by-30-ft room?

OpenStudy (anonymous):

Carpet area = 8*10 = 80 ft^2 new area = 24*30 = 720 ft^2 Divide new are by old and multiply by 350

OpenStudy (anonymous):

3150?

OpenStudy (anonymous):

yes that seems to be what i got

OpenStudy (anonymous):

thank you! :)

OpenStudy (anonymous):

no problem :D